OpenStudy (anonymous):

How do I write an equivalent expression using exponents with this question: (^3 sq. root of x^2)(^4 sq. root of x^3)?

OpenStudy (saifoo.khan):

Can you please ues the draw tool to write it clearly?

OpenStudy (saifoo.khan):

\[\Huge (\sqrt[3]{x^2})(\sqrt[4]{x^3})\] is it like that?

OpenStudy (anonymous):

yes :P

OpenStudy (saifoo.khan):

this can be written as: \[\Huge x^\frac23 \times x^\frac34\]

OpenStudy (saifoo.khan):

\[\huge x^{\frac23+\frac34}\]

OpenStudy (anonymous):

is it the same as x 17/12?

OpenStudy (anonymous):

because that's what the answer key has.. /:

OpenStudy (saifoo.khan):

Righto.

OpenStudy (saifoo.khan):

Solve it by taking LCM.

OpenStudy (saifoo.khan):

\[\huge x^\frac{17}{12}\]